OpenStudy (anonymous):

How would I change the equation into slope-intercept form? -3x + 2y = 8

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

-3x+2y=8 2y=3x+8 y=(3/2)x+4

OpenStudy (anonymous):

Okay, but how did you do that? Can you show me?

OpenStudy (anonymous):

and the (3/2) that wouldn't be negative?

OpenStudy (phi):

you start with -3x + 2y = 8 you want slope-intercept form ( you should remember this means y = mx + b) looking at y = mx +b as the pattern, the first thing you notice is y is by itself on the left side. A good step is to add +3x to both sides (because you know -3x + 3x =0 will make the -3x "go away" on the left side) -3x + 3x + 2y = +3x +8

OpenStudy (phi):

-3x + 3x + 2y = +3x +8 on the left side you have 3 x's take away 3 x's. That leaves 0 x's this simplifies to 2y = 3x + 8 now divide both sides by 2 y = (3x + 8)/2 or y = (3/2)x + 4

OpenStudy (anonymous):

this is how its done. -3x+2y=8 2y=3x+8 y=(3/2)x+4. u start with ur equation, then u bring the 3x to the other side and then u get rid of the "2" in the y

OpenStudy (anonymous):

so I add 3x to 2y and 3x?

OpenStudy (anonymous):

you add 3x to both sides of the equation yes

OpenStudy (anonymous):

Okay, Thank You (: